OpenStudy (anonymous):

what is the common ratio of -3,-12,-48,-192?

OpenStudy (solomonzelman):

common ratio, is the number by which you multiply each time to obtain the next number in the sequence.

OpenStudy (anonymous):

Looks like they are getting multiplied by 4 -3*4 =-12 -12*-4=-48 and so on.

OpenStudy (anonymous):

So the common ration is 4, @Yup11